# A Time Series Analysis of Assertions in the Linux Kernel

### v1

@@ 5. Conclusion
The handling of errors in the kernel space would also deserve more research. Indeed, whether and how well an operating system kernel handles bugs and fatal errors is a classical research question (Chou01, ). While there is a whole research branch on fault-tolerant operating systems and their kernels, it would be worthwhile to study also the error recovery strategies used in existing operating system kernels. This topic could be studied also empirically by focusing on frameworks and interfaces as well as programming patterns and idioms used for error handling in the kernel space. A comparative setup for empirical research could be constructed by again concentrating on the subsystems of the Linux kernel or enlarging the scope to other open source software operating system kernels. An interesting research path would open also by focusing on the potential differences between error handling with the C and Rust programming languages. Another option would be to continue with the themes noted in Section 2. For instance, the existing knowledge seems limited with respect to the alleged benefits of assertions for documenting code and helping debugging in the kernel context.

### v2

@@ 5 Conclusion
The handling of errors in the kernel space would also deserve more research. Indeed, whether, how, and how well an operating system kernel handles bugs and fatal errors are classical research questions [6, 9]. While there is a whole research branch on fault-tolerant operating systems and their kernels, it would be worthwhile to study also the error recovery strategies used in existing operating system kernels. The topic could be studied also empirically by focusing on frameworks and interfaces as well as programming patterns and idioms used for error handling in the kernel space. A comparative setup for empirical research could be constructed by again concentrating on the subsystems of the Linux kernel or enlarging the scope to other open source software operating system kernels. An interesting research path would open also by focusing on the potential differences between error handling with the C and Rust programming languages. Another option would be to continue with the themes noted in Section 2. For instance, the existing knowledge seems limited with respect to the alleged benefits of assertions for documenting code and helping debugging in the kernel context. Answers to these and related questions might help at reaching the goal of moving toward software assertion use in the Linux kernel.